# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ 5.5.3 Types of models
As for the proposed models, 35 out of 66 belong to the pure convolutional GNN paradigm, 12 to hybrid convolutional-attention architectures, followed by 14 attentional GNNs, and 1 recurrent GNN. Two papers use both convolutional and attentional models, and other two (Refs. [art_438, art_451]) do not fall into a specific class. Regarding the definition of the graph structure, most of the papers discuss it in detail. Half of the papers published before 2024 are based on a pre-defined adjacency matrix, while the majority of papers of 2024 propose models that learn the adjacency matrix on their own, either from scratch or based on a pre-initialization (e.g., Refs. [art_108, art_444]). Here, more than in any other thematic group, there are multiple ways to define the adjacency matrix. This can be made by taking into account the structure of the road network, the actual flow of people, or a combination of different sources of information. More in detail, a first group of papers uses road connectivity (e.g., Refs. [art_104, art_120, art_123, art_135, art_431]); second group uses spatial distance (e.g., Refs. [art_93, art_96, art_99, art_110, art_128, art_132, art_434, art_454]); a third group incorporates extra information about the so-called points of interest (such as Refs. [art_94, art_96]); a fourth group considers some similarity measures between the time series (e.g., Refs. [art_97, art_117, art_122, art_123, art_435, art_445]), such as the correlation coefficient or cosine similarity; a fifth group directly uses the number of people traveling on the road at a given time step (e.g., Refs. [art_94, art_102, art_129, art_133]), or a combination of different aspects (e.g., Refs. [art_420, art_444, art_448, art_449]).
The two most common loss functions used to train the models are the MAE and the MSE. Few papers consider a combination of loss functions, especially when dealing with a model that has to learn the graph structure by itself (Refs. [art_126, art_130, art_446, art_448]). Not all papers specify the language or library used for writing the code. Among those that do that, Python is the most common language, with the majority using PyTorch, whereas only a few researchers use TensorFlow. In addition, a few papers provide a link to the source code for the proposed model. The papers that provide such links are listed in Tab. LABEL:t:link_codici_traffico.
@@
3-Dimensional Graph Convolution Network (3DGCN) [art_94] https://github.com/FIBLAB/3D-DGCN
Adaptive generalized PageRank graph neural network (AGP-GNN) [art_116] https://github.com/guoxiaoyuatbjtu/agp-gnn
Adaptive multigraph convolutional networks (AMGCN) [art_444] https://github.com/hfimmortal/AMGCN
Automated dilated spatio-temporal synchronous graph network (Auto-DSTSGN) [art_108] https://github.com/jinguangyin/Auto-DSTSGN
BigST [art_447] https://github.com/usail-hkust/BigST
Dynamic multi-view graph neural network for citywide traffic inference (CTVI+) [art_97] https://github.com/dsj96/TKDD
Dynamic Graph Convolutional Recurrent Network (DGCRN) [art_100] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
Progressive Graph Convolutional Network (PGCN) [art_441] https://github.com/yuyolshin/PGCN
Scale-aware neural architecture search framework for MTS forecasting (SNAS4MTF) [art_408] https://github.com/shangzongjiang/SNAS4MTF
Spatio-temporal Causal Graph Attention Network (STCGAT) [art_125] https://github.com/zsqZZU/STCGAT/tree/v1.0.0
Spatial-Temporal Graph Convolutional Neural Network with LSTM layers (STGCN-LSTM) [3] https://github.com/ant-agafonov/stgcn-lstm
Spatio-Temporal Graph Mixformer (STGM) [art_110] https://github.com/mouradost/stgm
Space-Time Separable Graph Convolutional Network (STSGCN) [art_443] https://github.com/FraLuca/STSGCN
Two-level Resolution Neural Network (TwoResNet) [art_430] https://github.com/semink/TwoResNet
Wavelet dynamic spatiotemporal aware Graph Neural Network (W-DSTAGNN) [art_419] https://github.com/yash-jakhmola/w-dstagnn
@@ 5.5.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_93], [art_94], [art_96], [art_98], [art_100], [art_103], [art_106], [art_109], [art_116], [art_123], [art_126], [art_129], [art_132], [art_135], [art_406], [art_408], [art_419], [art_421], [art_424], [art_430], [art_433], [art_438], [art_444], [art_446], [art_448], [art_449], [art_451], [art_452], [art_453]
Historical average (HA) [art_93], [art_94], [art_96], [art_98], [art_100], [art_102], [art_103], [art_109], [art_110], [art_120], [art_123], [art_125], [art_131], [art_132], [art_135], [art_406], [art_408], [art_420], [art_424], [art_433], [art_435], [art_438], [art_441], [art_443], [art_446], [art_447], [art_449], [art_451], [art_452], [art_453]
Vector autoregression (VAR) [art_99], [art_100], [art_126], [art_131], [art_132], [art_406], [art_419], [art_424], [art_435], [art_443], [art_445], [art_447], [art_453]
Non-GNN machine learning methods Convolutional long-short term memory network (ConvLSTM) [ConvLSTM] [art_98], [art_446]
Deep multi-view spatial-temporal network (DMVST-Net) [DMVST-Net] [art_96], [art_451]
Feed-forward neural network (FNN) [art_93], [art_97], [art_100], [art_105], [art_114], [art_120], [art_121], [art_124], [art_126], [art_408], [art_420], [art_431], [art_438]
Gated recurrent unit (GRU) [art_98], [art_99], [art_106], [art_121], [art_123], [art_124], [3], [art_131], [art_132], [art_135], [art_425], [art_452]
Long-short term memory network (LSTM) [art_93], [art_96], [art_98], [art_99], [art_100], [art_106], [art_108], [art_109], [art_113], [art_115], [art_116], [art_120], [art_121], [art_122], [art_125], [art_126], [3], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_424], [art_425], [art_430], [art_433], [art_435], [art_438], [art_441], [art_442], [art_443], [art_444], [art_445], [art_449], [art_453]
Long-short term time series network (LSTNet) [LSTNet] [art_102], [art_110]
Seq2Seq [Seq2seq] [art_105], [art_124], [art_451]
@@
Spatial and temporal normalization for multivariate time series forecasting (ST-Norm) [ST-Norm] [art_116], [art_117], [art_406], [art_422]
Spatial-temporal identity (STID) [STID] [art_116], [art_422], [art_433], [art_434]
Support vector regression (SVR) [art_99], [art_100], [art_103], [art_109], [art_119], [art_123], [art_129], [art_130], [art_135], [art_408], [art_419], [art_421], [art_435], [art_444], [art_445], [art_453]
Temporal convolutional network (TCN) [TCN] [art_446], [art_453]
@@
Temporal 2D-variation modeling for general time series analysis (TimesNet) [TimesNet] [art_434], [art_435]
Transformer [attention] [art_424], [art_454]
Extreme gradient boosting (XGBoost) [XGBoost] [art_97], [art_114]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_98], [art_100], [art_110], [art_117], [art_124], [art_125], [art_130], [art_131], [art_406], [art_419], [art_435], [art_441], [art_447], [art_448], [art_452], [art_453], [art_454]
@@
Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_93], [art_98], [art_99], [art_100], [art_108], [art_110], [art_113], [art_115], [art_117], [art_121], [art_125], [art_126], [3], [art_130], [art_131], [art_132], [art_406], [art_421], [art_424], [art_427], [art_432], [art_433], [art_435], [art_442], [art_445], [art_446], [art_447], [art_448], [art_449], [art_453], [art_454]
@@
Attention based Spatial-Temporal Graph Neural Network (ASTGNN) [ASTGNN] [art_422], [art_435], [art_446]
Automated spatio-temporal graph (AutoSTG) [AutoSTG] [art_108], [art_408]
Dual-stage attention-based recurrent neural network (DA-RNN) [DA-RNN] [art_96], [art_105]
@@
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_93], [art_94], [art_96], [art_100], [art_103], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_122], [art_124], [art_125], [art_126], [art_128], [art_129], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_422], [art_424], [art_425], [art_427], [art_430], [art_432], [art_433], [art_435], [art_438], [art_441], [art_442], [art_444], [art_445], [art_447], [art_448], [art_451], [art_452], [art_453]
Dynamic graph convolutional recurrent network (DGCRN) [art_100] [art_110], [art_406], [art_444]
Dynamic and Multi-faceted SpatioTemporal Graph Convolution Network (DMSTGCN) [DMSTGCN] [art_441], [art_444]
Dynamic spatial-temporal aware graph neural network (DSTAGNN) [DSTAGNN] [art_125], [art_419], [art_424], [art_447], [art_448]
Fully connected gated graph architecture (FC-GAGA) [FC-GAGA] [art_116], [art_408], [art_427]
GeoMAN [GeoMAN] [art_96], [art_132]
Graph convolutional network (GCN) [art_104], [art_135], [art_431]
Graph multi-attention network (GMAN) [GMAN] [art_100], [art_109], [art_110], [art_116], [art_122], [art_128], [art_408], [art_419], [art_422], [art_427], [art_430], [art_444], [art_451]
Graph WaveNet (GWN) [GWN] [art_93], [art_100], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_122], [art_124], [art_128], [art_406], [art_408], [art_419], [art_420], [art_422], [art_424], [art_430], [art_432], [art_433], [art_434], [art_435], [art_441], [art_444], [art_447], [art_448], [art_449], [art_452], [art_453], [art_454]
Discrete graph structure learning for time series (GTS) [GTS] [art_420], [art_433]
Long short-term graph convolutional networks (LSGCN) [LSGCN] [art_99], [art_432], [art_446], [art_453]
Multi-range attentive bicomponent graph convolutional network (MRA-BGCN) [MRA-BGCN] [art_109], [art_116], [art_408]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_100], [art_109], [art_110], [art_113], [art_116], [art_117], [art_122], [art_124], [art_128], [art_406], [art_408], [art_420], [art_422], [art_444], [art_445]
Multi-view graph convolutional network (MVGCN) [MVGCN] [art_94], [art_96]
Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [StemGNN] [art_406], [art_422]
Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_98], [art_108], [art_115], [art_117], [art_125], [art_130], [art_406], [art_435], [art_445], [art_452], [art_453]
Spatial-temporal graph to sequence model (STG2Seq) [STG2Seq] [art_99], [art_424]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_93], [art_94], [art_96], [art_99], [art_100], [art_106], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_121], [art_122], [art_124], [art_126], [3], [art_128], [art_129], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_422], [art_424], [art_425], [art_433], [art_435], [art_438], [art_442], [art_443], [art_444], [art_445], [art_448], [art_449], [art_451], [art_452], [art_453], [art_454]
@@
Spatial temporal graph neural network (STGNN) [STGNN] [art_94], [art_103]
Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_99], [art_108], [art_117], [art_125], [art_406], [art_424], [art_447], [art_453]
Spatio-temporal graph attention network (ST-GRAT) [ST-GRAT] [art_116], [art_444]
Deep-meta-learning based model (ST-MetaNet) [ST-MetaNet] [art_96], [art_100], [art_109], [art_116], [art_408], [art_451]
Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_98], [art_99], [art_100], [art_108], [art_115], [art_117], [art_125], [art_130], [art_131], [art_406], [art_419], [art_424], [art_435], [art_441], [art_442], [art_443], [art_453], [art_454]
Temporal graph convolutional network (T-GCN) [T-GCN] [art_103], [art_113], [art_121], [art_123], [art_421], [art_433]
@@
Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [Z-GCNETs] [art_125], [art_424], [art_441]
@@
As for the basic mathematical and statistical methods, two widely used benchmarks are the ARIMA model and the historical average model (HA), which uses the average of past observations as the forecast for future values. Despite their apparent simplicity, these benchmarks are sometimes shown to be very powerful, with an accuracy comparable to that of more complex models. As for the non-GNN machine learning models, LSTM networks are widely recognized as a type of recurrent neural network capable of handling long-term dependencies in traffic data. As for GNN-based models instead, the three most used models are the spatio-temporal graph convolutional network (STGCN) [STGCN], the diffusion convolutional recurrent neural network (DCRNN) [DCRNN], and the attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], all developed specifically for the traffic forecasting problem.
@@
The STGCN model proposed by Yu et al. in 2018 consists of two spatio-temporal convolutional blocks (with two temporal gated convolutional layers, one spatial graph convolutional layer in the middle, and residual connections) and a fully-connected output layer in the end. In its original implementation, it processes graphs a with weighted adjacency matrix based on spatial distance between sensors in the road network. The source code for the STGCN model is available at https://github.com/VeritasYin/STGCN_IJCAI-18. The DCRNN model proposed by Li et al. in 2017 aims to capture the spatial dependency by using bidirectional random walks on the graph (whose adjacency matrix is based on node proximity), and to capture the temporal dependencies by using an encoder-decoder architecture with scheduled sampling. The code can be accessed at https://github.com/liyaguang/DCRNN. The ASTGCN model proposed by Guo et al. in 2019 models at once recent, daily-periodic and weekly-periodic dependencies as three independent components, with a spatio-temporal attention mechanism for capturing the dynamic spatio-temporal correlations and a spatio-temporal convolution which simultaneously employs graph convolutions to capture the spatial patterns. The source code for the ASTGCN model is available at https://github.com/Davidham3/ASTGCN-2019-mxnet. For a more comprehensive list of source codes for the most dated benchmark GNN models in the traffic domain see Ref. [review_traffico_jiang].

### v4

@@ 5.5.3 Types of models
As for the proposed models, 53 out of 102 belong to the pure convolutional GNN paradigm, followed by 21 hybrid convolutional-attention architectures, 20 attentional GNNs, and 1 recurrent GNN. Two papers use both convolutional and attentional models, and other five (Refs. [406, 388, 84, 370, 411]) do not fall into a specific class. Regarding the definition of the graph structure, most of the papers discuss it in detail. Almost half of the papers published before 2024 are based on a pre-defined adjacency matrix, while the majority of papers since 2024 propose models that learn the adjacency matrix on their own, either from scratch or based on a pre-initialization (e.g., Refs. [173, 208]). Here, more than in any other thematic group, there are multiple ways to define the adjacency matrix. This can be made by taking into account the structure of the road network, the actual flow of people, or a combination of different sources of information. More in detail, a first group of papers uses road connectivity (e.g., Refs. [489, 490, 169, 330, 134, 80]); second group uses spatial distance (e.g., Refs. [274, 211, 218, 200, 22, 27, 420, 193]); a third group incorporates extra information about the so-called points of interest (such as Refs. [428, 211]); a fourth group considers some similarity measures between the time series (e.g., Refs. [62, 111, 181, 510, 250]), such as the correlation coefficient or cosine similarity; a fifth group directly uses the number of people traveling on the road at a given time step (e.g., Refs. [428, 402, 25, 71]), or a combination of different aspects (e.g., Refs. [159, 208]). For a more detailed overview of graph construction techniques in traffic flow prediction models, the reader is referred to Ref. [78].
The two most common loss functions used to train the models are the MAE and the MSE. Few papers consider a combination of loss functions, especially when dealing with a model that has to learn the graph structure by itself (Refs. [39, 112, 89, 341, 276]). Not all papers specify the language or library used for writing the code. Among those that do that, Python is the most common language, with the majority using PyTorch, whereas only a few researchers use TensorFlow. In addition, a few papers provide a link to the source code for the proposed model. The papers that provide such links are listed in Tab. LABEL:t:link_codici_traffico.
@@
3-Dimensional graph convolution network (3DGCN) [428] https://github.com/FIBLAB/3D-DGCN
Adaptive graph convolutional recurrent network (AGCRN) [8] https://github.com/LeiBAI/AGCRN
Adaptive generalized PageRank graph neural network (AGP-GNN) [120] https://github.com/guoxiaoyuatbjtu/agp-gnn
Adaptive multigraph convolutional networks (AMGCN) [208] https://github.com/hfimmortal/AMGCN
Asynchronous spatio-temporal graph convolutional network (ASeer) [493] https://github.com/usail-hkust/ASeer
Automated dilated spatio-temporal synchronous graph network (Auto-DSTSGN) [173] https://github.com/jinguangyin/Auto-DSTSGN
BigST [126] https://github.com/usail-hkust/BigST
Dynamic multi-view graph neural network for citywide traffic inference (CTVI+) [62] https://github.com/dsj96/TKDD
Decoupled dynamic spatial-temporal graph neural network (D2STGNN) [342] https://github.com/zezhishao/D2STGNN
Domain adversarial spatial-temporal network (DastNet) [370] https://github.com/YihongT/DASTNet
Dynamic graph convolutional recurrent network (DGCRN) [212] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
Deep Kalman filtering network (DKFN) [33] https://github.com/Fanglanc/DKFN
Dynamic and multi-faceted spatio-temporal graph convolution network (DMSTGCN) [127] https://github.com/liangzhehan/DMSTGCN
Dynamic spatial-temporal aware graph neural network (DSTAGNN) [203] https://github.com/SYLan2019/DSTAGNN
Graph bootstrap (Gboot) [95] https://github.com/wangzz-yyzz/Gboot
Hierarchical information enhanced spatio-temporal (HIEST) [281] https://github.com/VAN-QIAN/CIKM23-HIEST
Multi-step dependency relation (MSDR) [259] https://github.com/dcliu99/MSDR
Multi-range spatial-temporal transformer network (MultiSPANS) [530] https://github.com/SELGroup/MultiSPANS
Progressive graph convolutional network (PGCN) [348] https://github.com/yuyolshin/PGCN
Scalable adaptive graph diffusion forecasting network (SAGDFN) [171] https://github.com/JIANGYUE61610306/SAGDFN
Scale-aware neural architecture search framework for MTS forecasting (SNAS4MTF) [32] https://github.com/shangzongjiang/SNAS4MTF
Spatio-temporal adaptive gated graph convolution network (STAG-GCN) [275] https://github.com/RobinLu1209/STAG-GCN
Spatial-temporal convolutional graph attention network (ST-CGA) [496] https://github.com/jbdj-star/cga
Spatio-temporal causal graph attention network (STCGAT) [511] https://github.com/zsqZZU/STCGAT/tree/v1.0.0
STGNN enhanced by a scalable time series pre-training model (STEP) [341] https://github.com/zezhishao/STEP
Spatial-temporal graph ordinary differential equation networks (STGODE) [84] https://github.com/square-coder/STGODE
Spatio-temporal graph mixformer (STGM) [200] https://github.com/mouradost/stgm
Spatio-temporal graph prompting (STGP) [144] https://github.com/hjf1997/STGP
Space-time separable graph convolutional network (STSGCN) [419] https://github.com/FraLuca/STSGCN
Spatio-temporal network with discrete wavelet transform (STWave) [81] https://github.com/LMissher/STWave
Sparse unstructured spatio temporal reconstruction (SUSTeR) [411] https://github.com/ywoelker/SUSTeR
Transferable graphs for traffics (TransGTR) [178] https://github.com/KL4805/TransGTR/
Two-level resolution neural network (TwoResNet) [199] https://github.com/semink/TwoResNet
Wavelet dynamic spatiotemporal aware graph neural network (W-DSTAGNN) [163] https://github.com/yash-jakhmola/w-dstagnn
@@ 5.5.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [274], [428], [211], [220], [212], [263], [262], [217], [120], [39], [25], [27], [169], [227], [32], [163], [385], [407], [199], [242], [406], [208], [89], [388], [82], [171], [127], [276], [393], [178], [81], [524], [84], [144], [85], [213], [222], [33], [320], [219], [496], [145], [275]
Historical average (HA) [274], [428], [211], [220], [212], [402], [263], [217], [200], [490], [511], [69], [27], [169], [227], [32], [159], [407], [242], [510], [406], [348], [419], [89], [126], [388], [82], [341], [127], [134], [493], [276], [135], [393], [81], [342], [524], [259], [8], [144], [222], [370], [33], [123], [320], [219], [275], [301]
$k$-nearest neighbors (KNN) [62], [144], [275]
Naive [493], [123], [301]
Vector autoregression (VAR) [218], [212], [39], [69], [27], [227], [163], [407], [510], [419], [250], [126], [82], [341], [171], [127], [393], [81], [342], [8], [530], [213], [461], [145], [301]
Non-GNN machine learning methods Adaptive recurrent neural network (AdaRNN) [75] [276], [144]
Convolutional long-short term memory network (ConvLSTM) [346] [220], [89], [123], [301]
Deep multi-view spatial-temporal network (DMVST-Net) [447] [211], [388], [496]
Feed-forward neural network (FNN) [274], [62], [212], [431], [245], [490], [506], [278], [39], [32], [159], [330], [406], [213], [214], [275], [301]
Gated recurrent unit (GRU) [220], [218], [262], [506], [278], [69], [27], [169], [312], [493], [8], [370], [219]
Long-short term memory network (LSTM) [274], [211], [220], [218], [212], [262], [173], [217], [198], [438], [120], [490], [506], [181], [511], [39], [112], [69], [27], [227], [32], [163], [159], [385], [407], [312], [199], [242], [510], [406], [348], [378], [419], [208], [250], [82], [341], [171], [134], [493], [135], [393], [81], [342], [524], [203], [259], [80], [530], [85], [213], [387], [222], [33], [219], [214], [461], [496], [145], [275], [301]
Long-short term time series network (LSTNet) [202] [402], [200]
Multitask deep-learning (MDL) [482] [388], [523]
Recurrent neural network (RNN) [388], [286], [135], [393]
Seq2Seq [365] [431], [278], [388]
Spatial-temporal dynamic network (STDN) [446] [388], [496]
Spatial-temporal identity (STID) [340] [120], [466], [242], [420], [524], [422], [387]
@@
Spatial and temporal normalization for multivariate time series forecasting (ST-Norm) [67] [120], [111], [227], [466]
Spatio-temporal residual networks (ST-ResNet) [480] [388], [496]
Support vector regression (SVR) [218], [212], [263], [217], [251], [25], [112], [169], [32], [163], [385], [510], [208], [250], [82], [341], [171], [393], [81], [342], [530], [422], [213], [370], [33], [320], [461], [281], [496]
Temporal convolutional network (TCN) [9] [89], [82], [493], [81], [203], [80], [63]
@@
Temporal 2D-variation modeling for general time series analysis (TimesNet) [413] [420], [510]
Transformer [375] [407], [193], [523]
Extreme gradient boosting (XGBoost) [43] [62], [245], [127]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [8] [220], [212], [200], [111], [278], [511], [112], [69], [227], [163], [510], [348], [126], [82], [193], [171], [81], [524], [203], [80], [95], [213], [63], [370], [33], [214], [145]
@@
Attention-based spatial-temporal graph convolutional network (ASTGCN) [116] [274], [220], [218], [212], [173], [200], [198], [438], [111], [506], [511], [39], [112], [69], [27], [227], [385], [407], [13], [352], [242], [510], [378], [250], [89], [126], [82], [193], [341], [171], [81], [342], [524], [203], [84], [259], [8], [80], [95], [530], [422], [85], [213], [63], [33], [123], [320], [214], [461], [145], [275]
@@
Attention based spatial-temporal graph neural network (ASTGNN) [117] [466], [510], [89]
Automated spatio-temporal graph (AutoSTG) [305] [173], [32], [213], [33], [214]
@@
Coupled layer-wise convolutional recurrent neural network (CCRNN) [449] [259], [422]
@@
Decoupled dynamic spatial-temporal graph neural network (D2STGNN) [342] [171], [411]
Dual-stage attention-based recurrent neural network (DA-RNN) [318] [211], [431]
Diffusion convolutional recurrent neural network (DCRNN) [237] [274], [428], [211], [212], [263], [173], [217], [200], [198], [438], [120], [111], [490], [181], [278], [511], [39], [22], [25], [112], [69], [27], [227], [32], [163], [159], [385], [466], [407], [312], [13], [199], [352], [242], [510], [406], [348], [378], [208], [250], [126], [388], [82], [341], [171], [127], [493], [393], [81], [342], [524], [203], [84], [122], [259], [8], [144], [80], [95], [530], [422], [85], [213], [387], [222], [370], [33], [123], [214], [461], [281], [145], [275], [308], [301]
Deep spatial-temporal network (DeepST) [481] [388], [496]
Dynamic graph convolutional recurrent network (DGCRN) [212] [200], [227], [208], [342]
Dynamic and multi-faceted spatiotemporal graph convolution network (DMSTGCN) [127] [348], [208]
Dynamic spatial-temporal aware graph neural network (DSTAGNN) [203] [511], [163], [407], [126], [144], [80], [95], [63], [461], [145]
Fully connected gated graph architecture (FC-GAGA) [470] [120], [32], [13]
Gated attention networks for learning on large and spatiotemporal graphs (GaAN) [479] [393], [308]
Graph convolutional network (GCN) [489], [169], [330], [135], [370]
GeoMAN [246] [211], [27]
Graph multi-attention network (GMAN) [513] [212], [217], [200], [120], [181], [22], [32], [163], [466], [13], [199], [208], [388], [341], [171], [127], [342], [524], [530], [85], [213], [222], [33], [214], [461], [281], [145], [308]
Discrete graph structure learning for time series (GTS) [338] [159], [242], [341], [171], [178], [524], [387], [281]
Graph WaveNet (GWN) [425] [274], [212], [173], [217], [200], [198], [438], [120], [111], [181], [278], [22], [227], [32], [163], [159], [466], [407], [199], [352], [242], [420], [510], [348], [208], [126], [82], [193], [341], [171], [127], [493], [178], [81], [342], [524], [84], [122], [259], [144], [95], [530], [422], [85], [213], [387], [523], [222], [33], [214], [461], [281], [145], [308], [301]
Hierarchical graph convolution network (HGCN) [115] [385], [281]
Long short-term graph convolutional networks (LSGCN) [154] [218], [352], [89], [82], [81]
Multi-graph convolutional networks (MGCN) [30] [245], [134]
Multi-range attentive bicomponent graph convolutional network (MRA-BGCN) [44] [217], [120], [32], [222]
Multivariate time series forecasting with graph neural network (MTGNN) [424] [212], [217], [200], [198], [120], [111], [181], [278], [22], [227], [32], [159], [466], [208], [250], [341], [171], [127], [342], [524], [530], [213], [222], [33], [214], [281]
Multi-view graph convolutional network (MVGCN) [359] [428], [211]
Propagation delay-aware dynamic long-range transformer (PDFormer) [165] [242], [493], [95], [461]
a Spatio-temporal adaptive embedding transformer (STAEformer) [260] [493], [63]
Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [26] [227], [466], [387], [123]
STGNN enhanced by a scalable time series Pre-training model (STEP) [341] [171], [144], [387]
Spatial-temporal fusion graph neural network (STFGNN) [228] [220], [173], [438], [111], [511], [112], [227], [510], [250], [82], [81], [203], [259], [80], [95], [213], [33], [214], [145]
Spatial-temporal graph to sequence model (STG2Seq) [7] [218], [407], [259], [80], [422], [523]
Spatio-temporal graph convolutional network (STGCN) [457] [274], [428], [211], [218], [212], [262], [173], [217], [200], [198], [438], [120], [111], [490], [506], [181], [278], [39], [22], [25], [112], [69], [27], [227], [32], [163], [159], [385], [466], [407], [312], [242], [510], [406], [378], [419], [208], [250], [388], [82], [193], [171], [81], [203], [84], [122], [259], [8], [80], [95], [530], [85], [213], [387], [222], [370], [33], [214], [461], [281], [496], [145], [275], [308], [301], [411]
Spatio-temporal graph few shot learning (ST-GFSL) [276] [178], [144]
Spatial temporal graph neural network (STGNN) [393] [428], [263], [393]
Spatial-temporal graph ODE networks (STGODE) [84] [218], [173], [111], [511], [227], [407], [126], [82], [81], [203], [122], [80], [95], [422], [213], [370], [33]
Spatio-temporal graph attention network (ST-GRAT) [308] [120], [208]
Deep-meta-learning based model (ST-MetaNet) [306] [211], [212], [217], [120], [32], [388], [524], [213], [222], [33], [214], [496]
Spatiotemporal multi-graph convolution network (ST-MGCN) [101] [388], [496]
Spatio-temporal synchronous graph convolutional network (STSGCN) [353] [220], [218], [212], [173], [438], [111], [511], [112], [69], [227], [163], [407], [510], [348], [378], [419], [82], [193], [341], [171], [81], [342], [524], [203], [84], [259], [8], [80], [95], [213], [33], [214], [461], [145]
Temporal graph convolutional network (T-GCN) [508] [263], [198], [506], [385], [242], [393], [530], [422], [370], [320], [281]
@@
Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [50] [511], [407], [348], [203], [80], [145]
@@
As for the basic mathematical and statistical methods, two widely used benchmarks are the ARIMA model and the historical average model (HA), which uses the average of past observations as the forecast for future values. Despite their apparent simplicity, these benchmarks are sometimes shown to be very powerful, with an accuracy comparable to that of more complex models. As for the non-GNN machine learning models, LSTM networks are widely recognized as a type of recurrent neural network capable of handling long-term dependencies in traffic data. As for GNN-based models instead, the most used models are the spatio-temporal graph convolutional network (STGCN) [457], the diffusion convolutional recurrent neural network (DCRNN) [237], Graph WaveNet (GWN) [425], and the attention-based spatial-temporal graph convolutional network (ASTGCN) [116], all developed specifically for the traffic forecasting problem.
@@
The STGCN model proposed by Yu et al. in 2018 consists of two spatio-temporal convolutional blocks (with two temporal gated convolutional layers, one spatial graph convolutional layer in the middle, and residual connections) and a fully-connected output layer in the end. In its original implementation, it processes graphs a with weighted adjacency matrix based on spatial distance between sensors in the road network. The source code for the STGCN model is available at https://github.com/VeritasYin/STGCN_IJCAI-18. The DCRNN model proposed by Li et al. in 2017 aims to capture the spatial dependency by using bidirectional random walks on the graph (whose adjacency matrix is based on node proximity), and to capture the temporal dependencies by using an encoder-decoder architecture with scheduled sampling. The code can be accessed at https://github.com/liyaguang/DCRNN. The GWN model was introduced by Wu et al. in 2021 as a novel approach to spatio-temporal forecasting, particularly in traffic tasks. It is known for its ability to learn an adaptive dependency matrix from node embeddings. The model combines graph convolutional and temporal convolutional layers to effectively capture spatio-temporal dependencies. The source code is available at https://github.com/nnzhan/Graph-WaveNet. The ASTGCN model proposed by Guo et al. in 2019 models at once recent, daily-periodic and weekly-periodic dependencies as three independent components, with a spatio-temporal attention mechanism for capturing the dynamic spatio-temporal correlations and a spatio-temporal convolution which simultaneously employs graph convolutions to capture the spatial patterns. The source code for the ASTGCN model is available at https://github.com/Davidham3/ASTGCN-2019-mxnet. For a more comprehensive list of source codes for the most dated benchmark GNN models in the traffic domain see Ref. [168].